# InternVideo2: Scaling Foundation Models for Multimodal Video Understanding

### v2

@@ 1 Introduction
Learning transferrable spatiotemporal representations is a critical research area in computer vision, holding diverse applications across domains such as video search Gabeur et al. (2020), game control Bruce et al. (2024), robotic learning Driess et al. (2023), self-driving Zablocki et al. (2022), and scientific studies Team et al. (2023). Recently, the advancement of Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023a); Touvron et al. (2023a, b) and their multimodal variations (MLLMs) OpenAI (2023b); Gong et al. (2023); Liu et al. (2023); Team et al. (2023) have had a profound impact on vision research and other disciplines. Embedding videos effectively into these large models and harnessing their capabilities to enhance video understanding performance has emerged as pivotal tasks Li et al. (2023c); Maaz et al. (2023b).
@@ 2 Related Work
Video Foundation Models. Studies on learning video foundation models becomes increasingly crucial considering its wide applications Li and Wang (2020); Xu et al. (2021); Li et al. (2023e); Wang et al. (2022); Zhao et al. (2024); Wang et al. (2023c); Yan et al. (2022); Feichtenhofer et al. (2022); Wang et al. (2023a); Tong et al. (2022); Wang et al. (2023b). Typical methods in building video foundation models (ViFM) include video-text contrastive learning Li and Wang (2020); Xu et al. (2021); Wang et al. (2022), masked video modeling Tong et al. (2022); Wang et al. (2023b, 2022); Fu et al. (2021), and next token prediction Alayrac et al. (2022); Sun et al. (2023c, a). Specifically, All-in-one Wang et al. (2023a) utilized a single backbone with unified multiple pretraining objectives. On the other hand, UMT Li et al. (2023e) combined masked modeling with video-text contrastive learning, demonstrating strong performance in both action recognition and video-language tasks. Another approach is mPLUG-2 Xu et al. (2023), which introduced a new design for modulating different modalities. It shared a common module across modalities to enhance relations while incorporating modality-specific modules for discrimination. In addition to video-text pretraining, researchers have also explored the use of audio information in videos to improve performance. MERLOT Reserve Zellers et al. (2022) learned video representations using a large-scale dataset of video-speech-transcript pairs. VALOR Chen et al. (2023b) employed independent encoders for video, audio, and text and trains a joint visual-audio-text representation. VAST Chen et al. (2024b) constructed an audio-visual-speech dataset and develops a multimodal backbone that excels in video-audio-related tasks. VideoPrism Zhao et al. (2024) combined video-text contrastive learning and video token reconstruction on a combination of public and proprietary videos, achieving leading results across various video tasks.
Multimodal Large Language Models. With advances in large language models (LLMs) Devlin et al. (2018); Raffel et al. (2020); Brown et al. (2020), their multimodal versions (MLLMs) is becoming popular as it can handle open-world tasks. Seminal works like Flamingo Alayrac et al. (2022) showed outstanding zero/few-shots performances over a range of multimodal tasks Goyal et al. (2017b); Plummer et al. (2015); Xu et al. (2016); Marino et al. (2019). Public MLLMs Zhu et al. (2023b); Liu et al. (2023); Gong et al. (2023) such as LLaVA Liu et al. (2023) and InstructBLIP Dai et al. (2023) proposed to use visual instruction-tuning data to improve the visual dialogue ability. Some video-centric MLLMs have been proposed, such as VideoChat Li et al. (2023c), VideoChatGPT Maaz et al. (2023b) and Valley Luo et al. (2023), by using instruction data to connect video encoders to LLMs for open-world video understanding.
@@ 3.2 Stage 2: Aligning Video to Audio-Speech-Text
We exploit the correspondence between video and audio, speech, and text to encourage InternVideo2 to learn more semantics. In practice, InternVideo2 has a huge video encoder, and its employed audio and text encoders are relatively lightweight. The used audio encoder is a 12-layer transformer initialized with BEATs Chen et al. (2023a) (90M). It takes in audio features, which are 64-dimensional log Mel filterbank spectrograms using a 25ms Hamming window, transformed from 10-second-long clips, padding with zeros. For the text and speech encoders, we initialize the text encoder and multimodal decoder using BERT-Large Devlin et al. (2018). Specifically, we utilize the initial 19 layers of BERT-Large as the text encoder, with the subsequent 5 layers equipped with cross-attention layers serving as the multimodal decoder.
@@ 4.3 Instruction-Tuning Data for Video Dialogue
We employ a updated training version of MVBench Li et al. (2023d). Originally, it comprises 1.9M samples (both images and videos) from 34 distinct sources. We decrease the amount of caption data from WebVid and CoCo to 80k and 100k separately and add new data from the S-MiT to boost the diversity rather than quantity of the instruction dataset. In the additional HD training stage, we incorporate the videos with corresponding GPT-4 annotations from Chen et al. (2024a) into the training set. We further expand the training set by including datasets from PerceptionTestQA Patraucean et al. (2024), TVQA Lei et al. (2018), NTU-RGB-D Liu et al. (2020), and EgotaskQA Grauman et al. (2022), along with grounding datasets based on DiDeMo Anne Hendricks et al. (2017b) and COCO Lin et al. (2014). This training data encompasses key features of image and video understanding across crucial tasks, including 1) conversation, 2) caption, 3) visual question answer, 4) reasoning, and 5) classification.
@@ 5.1.1 Action Recognition
We test InternVideo2 on Kinetics (i.e., K400, 600 and 700 Carreira and Zisserman (2017); Carreira et al. (2018, 2019)), Moments in Time V1 (MiT) Monfort et al. (2020), Something-Something V2 (SSv2) Goyal et al. (2017a), UCF Soomro et al. (2012), HMDB Kuehne et al. (2011), Charades Gao et al. (2017), ActivityNet Heilbron et al. (2015) (ANet) and HACS Zhao et al. (2019). We evaluate in four settings: (a) end-to-end finetuning the whole backbone; (b) attentive probing is similar to linear pooling, but extra trains the attention pooling layer Yu et al. (2022). (c) linear probing which freezes the backbone and only trains the task head; and (d) zero-shot.
@@
Tab. 2 shows InternVideo2-6B obtains new state-of-the-art (SOTA) results on Kinetics (92.1%/91.9%/85.9% on K400/600/700, respectively), SthSthv2, MiT, ANet, and HACS with only 16 frames, while the previous SOTAs require larger resolution (224 vs. 576) or model ensemble. As for MiT in Table 2, InternVideo2-6B exceeds the previous SOTA, CoCa-g, by a significant margin of 2.2% (51.2% vs. 49.0%). Regarding the temporal-related actions in Table 2, our InternVideo2-6B also surpasses MVD Wang et al. (2023c) on SSv2 (77.5% vs. 77.3%). Moreover, our InternVideo2-6B showcases top performance on untrimmed video analysis, as indicated in Table 2, with 95.9% on ActivityNet and 97.0% on HACS. These results affirm our model’s superior capability for robustly identifying complex actions across varied scenes. Note “I" and “V" denotes images and videos, respectively. “IV-3B" means the total number of the used images and videos is 3B, while “I-3B" means using 3B images.
@@
EVA-CLIP-18B Sun et al. (2024) 16 79.4 - 79.4 - 72.2
VideoPrism-g Zhao et al. (2024) 16 76.4 85.4 -
@@ 5.1.2 Temporal Action Localization
We evaluate models on four temporal action localization (TAL) datasets: THUMOS14 Idrees et al. (2017), ActivityNet Krishna et al. (2017), HACS Segment Zhao et al. (2019) and FineAction Liu et al. (2022) in a feature-based manner with finetuning. We employ output of the 7-th layer from InternVideo2 for inputs as the corresponding features without fusing anything else. ActionFormer Zhang et al. (2022) is used as the detection head. We report mean Average Precision(mAP) under multiple tIoU as in Lin et al. (2019); Chen et al. (2022a); Zhang et al. (2022); Yang et al. (2023). In Table 7, InternVideo2-6B gets the highest mAP among all comparisons in all datasets, while InternVideo2-1B nearly surpass other methods except in THUMOS14. We find InternVideo2-6B almost consistently improves mAP with a notable margin from InternVideo2-1B except in FineAction. We suppose scaling model capacity without data refinement cannot nontrivially improve fine-grained discrimination abilities of models. Scaling detailed annotations in training may address this issue.
@@ 5.2.1 Video Retrieval
We evaluate the video retrieval on six popular benchmarks Heilbron et al. (2015); Xu et al. (2016); Rohrbach et al. (2015); Anne Hendricks et al. (2017a); Chen and Dolan (2011), as shown in Tab. 9 and 10. In eval, eight frames from the input videos are uniformly sampled. We report R@1 scores for both text-to-video (t2v) and video-to-text (v2t) tasks in Tab. 9 and 10. R@5 and R@10 are given in Supp.
Table 9 and 10 demonstrate that InternVideo2 outperforms other state-of-the-arts with a notable margin in both t2v and v2t of all used datasets no matter in zero-shot or finetuned settings, except for the v2t of MSR-VTT, where VideoPrism gives the best result. This shows the video-language semantic alignment of transferrity of InternVideo2.
@@ 5.2.2 Video Temporal Grounding
We evaluate InternVideo2 on two video temporal grounding (VTG) datasets: Qvhighlight Lei et al. (2021), and Charade-STA Gao et al. (2017). The eval setting and used features are the same as in TAL. We use CG-DETR Moon et al. (2023) as the grounding head. We report R1@0.3, R1@0.5, R1@0.7, and mAP for moment retrieval as in Lei et al. (2021); Moon et al. (2023); Lin et al. (2023). Highlight Detection is evaluated in terms of “Very Good” mAP and HiT@1. In Table 11(b), InternVideo2-1B and InternVideo2-6B bring gradual performance improvements compared to CLIP Radford et al. (2021a) and CLIP Radford et al. (2021a)+Slowfast Feichtenhofer et al. (2019). This suggests that a larger spatiotemporal model is more beneficial to short-term video semantic alignment capabilities.
@@ 5.3 Video-centric Dialogue and its Applications
Tab. 14 shows results on MVBench Li et al. (2023d), Egoschema Mangalam et al. (2023), and Perception Test Patraucean et al. (2023), by equipping VideoChat2 Li et al. (2023d) with our InternVideo2 encoder, compared with other MLLMs. Note VideoChat2HD-InternVideo2 outperforms other systems with a notable margin in Perception test and MVBench (F16 denotes the model is trained and tested with 16-frame inputs), except in Egoschema compared with proprietary commercial models (GPT4 & Gemini). Our method demonstrates superior short-term fine-grained video understanding compared to both open-source and commercial methods. Egoschema requires longer context harnessing, and we plan to further explore leveraging LLM for long multimodal sequence modeling. Considering these benchmarks not only involve perception but also reasoning, these results suggest InternVideo2 does embed knowledge to partially model the world. It also validates the importance of learning a transferrable video representation for current video-related MLLM. We also give several qualitative evaluations with popular GPT-4V and GeminiPro, including dialogues on action sequence (Figure 4), confused action (Figure 5), temporal order understanding (Figure 6), temporal event counting (Figure 7, unexpected action reasoning (Figure 8), and vision-language navigation (Figure 9).
@@ 5.4.2 Training Data and used Teachers in Stage 1
In Tab. 24, we examine the impact of distillation teachers and used dataset sizes on model performance. (a) Data Scale: Note that pretraining data scale for MAE should grow with the increasing model scale, otherwise the downstream performance would be saturated, such as K710 (0.66M videos) for ViT-L, K-Mash1.1M for ViT-1B, and K-Mash2M for ViT-6B. (b) Teacher: Tab. 24 reveals that the synergy between a multimodal teacher (e.g., CLIP Radford et al. (2021a)) and a motion-aware teacher (e.g., MAE Tong et al. (2022)) markedly boosts performance, especially on SthSthV2. It highlights the importance of strategic teacher model selection in the distillation process.
@@ 5.4.4 Training and Evaluation in Stage 3
As shown in Table 19, incorporating questions (i.e., the ‘q’ in ‘QA’) into QFormer during Stage 3 training, which was found useful in Dai et al. (2023), actually harms the out-of-domain performance of the scaled-up VideoLLM. The NextQA training data is already included in the training corpus, and this is the only benchmark where the question-injected version performs better. Therefore, we believe that adding questions to QFormer during the instruction tuning stage of the scaled VideoChat model leads to some degree of overfitting.
@@ 6 Conclusion and Discussion
We introduce a novel video foundation model called InternVideo2, which achieves state-of-the-art performance across various video and audio tasks. In InternVideo2, we combine masked video token reconstruction, video-audio-text contrastive learning, and next token prediction into a unified framework. Additionally, we create a new video-text dataset that incorporates video-audio-speech fused captions as descriptions. The dataset contains temporally segmented clips with semantic differences. These designs in InternVideo2 contribute to enhancing video understanding in both perception and reasoning tasks. Notably, InternVideo2 excels in video-related dialogue and long video understanding, demonstrating its effectiveness in various video-related tasks.
@@
Despite its achievements, InternVideo2 does not introduce novel training methods or architectural advances. Instead, it leverages existing learning techniques for scheme exploration while focusing on improving data processing to enhance spatiotemporal perception, semantic alignment, and basic knowledge embedding. Similarly to previous studies Li et al. (2023e); Ye et al. (2023), the model still grapples with limitations stemming from fixed input resolutions, sampling rates, and highly compressed tokens, which restrict its ability to express rich video information and capture fine-grained details.
@@
Although InternVideo2 has demonstrated leading performance in long video understanding and reasoning benchmarks, it cannot guarantee an implicit world model that ensures consistency in visual reasoning. The inherent constraints imposed by fixed input representations, coupled with the complexity of visual reasoning tasks, present challenges in achieving a comprehensive and consistent understanding of the visual world.
@@
We investigate the potential biases here. We focused on age, gender, and race distributions, as these are commonly recognized areas where bias can occur. We count keywords related to these categories in the used captions. Note that these synthetic captions may not fully reflect the truth of the corresponding videos, thereby creating a gap between our analysis and the actual reality. Here are the results of our analysis:

### v3

@@ 1 Introduction
Learning transferrable spatiotemporal representations is a critical research area in computer vision, holding diverse applications across domains such as video search Gabeur et al. (2020), game control Bruce et al. (2024), robotic learning Driess et al. (2023), self-driving Zablocki et al. (2022), and scientific studies Team et al. (2023). Recently, the advancement of Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023a); Touvron et al. (2023a, b) and their multimodal variations (MLLMs) OpenAI (2023b); Gong et al. (2023); Liu et al. (2023); Team et al. (2023) have had a profound impact on vision research and other disciplines. Embedding videos effectively into these large models and harnessing their capabilities to enhance video understanding performance has emerged as pivotal tasks Li et al. (2023c); Maaz et al. (2023).
@@ 2 Related Work
Video Foundation Models. Studies on learning video foundation models become increasingly crucial considering its wide applications Li and Wang (2020); Xu et al. (2021); Li et al. (2023e); Wang et al. (2022); Zhao et al. (2024); Wang et al. (2023c); Yan et al. (2022); Feichtenhofer et al. (2022); Wang et al. (2023a); Tong et al. (2022); Wang et al. (2023b). Typical methods in building video foundation models (ViFM) include video-text contrastive learning Li and Wang (2020); Xu et al. (2021); Wang et al. (2022), masked video modeling Tong et al. (2022); Wang et al. (2023b, 2022); Fu et al. (2021), and next token prediction Alayrac et al. (2022); Sun et al. (2023c, a). Specifically, All-in-one Wang et al. (2023a) utilized a single backbone with unified multiple pretraining objectives. On the other hand, UMT Li et al. (2023e) combined masked modeling with video-text contrastive learning, demonstrating strong performance in both action recognition and video-language tasks. Another approach is mPLUG-2 Xu et al. (2023), which introduced a new design for modulating different modalities. It shared a common module across modalities to enhance relations while incorporating modality-specific modules for discrimination. In addition to video-text pretraining, researchers have also explored the use of audio information in videos to improve performance. MERLOT Reserve Zellers et al. (2022) learned video representations using a large-scale dataset of video-speech-transcript pairs. VALOR Chen et al. (2023b) employed independent encoders for video, audio, and text and trains a joint visual-audio-text representation. VAST Chen et al. (2024b) constructed an audio-visual-speech dataset and develops a multimodal backbone that excels in video-audio-related tasks. VideoPrism Zhao et al. (2024) combined video-text contrastive learning and video token reconstruction on a combination of public and proprietary videos, achieving leading results across various video tasks.
Multimodal Large Language Models. With advances in large language models (LLMs) Devlin et al. (2018); Raffel et al. (2020); Brown et al. (2020), their multimodal versions (MLLMs) is becoming popular as it can handle open-world tasks. Seminal works like Flamingo Alayrac et al. (2022) showed outstanding zero/few-shots performances over a range of multimodal tasks Goyal et al. (2017b); Plummer et al. (2015); Xu et al. (2016); Marino et al. (2019). Public MLLMs Zhu et al. (2023b); Liu et al. (2023); Gong et al. (2023) such as LLaVA Liu et al. (2023) and InstructBLIP Dai et al. (2023) proposed to use visual instruction-tuning data to improve the visual dialogue ability. Some video-centric MLLMs have been proposed, such as VideoChat Li et al. (2023c), VideoChatGPT Maaz et al. (2023) and Valley Luo et al. (2023), by using instruction data to connect video encoders to LLMs for open-world video understanding.
@@ 3.2 Stage 2: Aligning Video to Audio-Speech-Text
We exploit the correspondence between video and audio, speech, and text to encourage InternVideo2 to learn more semantics. In practice, InternVideo2 has a huge video encoder, and its employed audio and text encoders are relatively lightweight. The used audio encoder is a 12-layer transformer initialized with BEATs Chen et al. (2023a) (90M). It takes as input 64-dimensional log Mel filterbank spectrograms, generated using a 25ms Hamming window, from 10-second-long clips (padded with zeros). For the text and speech encoders, we initialize the text encoder and multimodal decoder using BERT-Large Devlin et al. (2018). Specifically, we utilize the initial 19 layers of BERT-Large as the text encoder, with the subsequent 5 layers equipped with cross-attention layers serving as the multimodal decoder.
@@ 4.3 Instruction-Tuning Data for Video Dialogue
We employ a updated training version of MVBench Li et al. (2023d). Originally, it comprises 1.9M samples (both images and videos) from 34 distinct sources. We decrease the amount of caption data from WebVid and CoCo to 80k and 100k separately and add new data from the S-MiT to boost the diversity rather than quantity of the instruction dataset. In the additional HD training stage, we incorporate the videos with corresponding GPT-4 annotations from Chen et al. (2024a) into the training set. We further expand the training set by including datasets from PerceptionTestQA Patraucean et al. (2024), TVQA Lei et al. (2018), NTU-RGB-D Liu et al. (2020), and EgotaskQA Grauman et al. (2022), along with grounding datasets based on DiDeMo Anne Hendricks et al. (2017) and COCO Lin et al. (2014). This training data encompasses key features of image and video understanding across crucial tasks, including 1) conversation, 2) caption, 3) visual question answer, 4) reasoning, and 5) classification.
@@ 5.1.1 Action Recognition
We test InternVideo2 on Kinetics (i.e., K400, 600, and 700 Carreira and Zisserman (2017); Carreira et al. (2018, 2019)), Moments in Time V1 (MiT) Monfort et al. (2020), Something-Something V2 (SSv2) Goyal et al. (2017a), UCF Soomro et al. (2012), HMDB Kuehne et al. (2011), Charades Gao et al. (2017), ActivityNet Heilbron et al. (2015) (ANet) and HACS Zhao et al. (2019). We evaluate in four settings: (a) end-to-end finetuning the whole backbone; (b) attentive probing is similar to linear pooling, but extra trains the attention pooling layer Yu et al. (2022). (c) linear probing which freezes the backbone and only trains the task head; and (d) zero-shot.
@@
Tab. 2 shows InternVideo2-6B obtains new state-of-the-art (SOTA) results on Kinetics (92.1%/91.9%/85.9% on K400/600/700, respectively), SthSthv2, MiT, ANet, and HACS with only 16 frames, while the previous SOTAs require larger resolution (224 vs. 576) or model ensemble. As for MiT in Tab. 2, InternVideo2-6B exceeds the previous SOTA, CoCa-g, by a significant margin of 2.2% (51.2% vs. 49.0%). Regarding the temporal-related actions in Tab. 2, our InternVideo2-6B also surpasses MVD Wang et al. (2023c) on SSv2 (77.5% vs. 77.3%). Moreover, our InternVideo2-6B showcases top performance on untrimmed video analysis, as indicated in Tab. 2, with 95.9% on ActivityNet and 97.0% on HACS. These results affirm our model’s superior capability for robustly identifying complex actions across varied scenes. Note “I" and “V" denotes images and videos, respectively. “IV-3B" means the total number of the used images and videos is 3B, while “I-3B" means using 3B images.
@@
EVA-CLIP-18B Sun et al. (2024) 16 - 79.4 - 79.4 - 72.2
VideoPrism-g Zhao et al. (2024) 16 76.4 85.4 - - - -
@@ 5.1.2 Temporal Action Localization
We evaluate models on four temporal action localization (TAL) datasets: THUMOS14 Idrees et al. (2017), ActivityNet Krishna et al. (2017), HACS Segment Zhao et al. (2019) and FineAction Liu et al. (2022) in a feature-based manner with finetuning. We employ output of the 7-th layer from InternVideo2 for inputs as the corresponding features without fusing anything else. ActionFormer Anne Hendricks et al. (2017) is used as the detection head. We report mean Average Precision(mAP) under multiple tIoU as in Lin et al. (2019); Chen et al. (2022a); Anne Hendricks et al. (2017); Yang et al. (2023). In Table 7, InternVideo2-6B gets the highest mAP among all comparisons in all datasets, while InternVideo2-1B nearly surpass other methods except in THUMOS14. We find InternVideo2-6B almost consistently improves mAP with a notable margin from InternVideo2-1B except in FineAction. We suppose scaling model capacity without data refinement cannot nontrivially improve fine-grained discrimination abilities of models. Scaling detailed annotations in training may address this issue.
@@ 5.2.1 Video Retrieval
We evaluate the video retrieval on six popular benchmarks: MSR-VTT Xu et al. (2016), LSMDC Rohrbach et al. (2015), DiDeMo Anne Hendricks et al. (2017), MSVD Chen and Dolan (2011), ActivityNet (ANet) Heilbron et al. (2015), and VATEX Wang et al. (2019), as shown in Tab. 9 and 10. In evaluation, eight frames from the input videos are uniformly sampled. We report R@1 scores for both text-to-video (t2v) and video-to-text (v2t) tasks in Tab. 9 and 10. R@5 and R@10 are given in Supp.
Tab. 9 and 10 demonstrate that InternVideo2 outperforms other state-of-the-arts with a notable margin in both t2v and v2t of all used datasets no matter in zero-shot or finetuned settings, except for the v2t of MSR-VTT, where VideoPrism gives the best result. This shows the video-language semantic alignment of transferrity of InternVideo2.
@@ 5.2.2 Video Temporal Grounding
We evaluate InternVideo2 on two video temporal grounding (VTG) datasets: QVhighlight Lei et al. (2021), and Charade-STA Gao et al. (2017). The eval setting and used features are the same as in TAL. We use CG-DETR Moon et al. (2023) as the grounding head. We report R1@0.3, R1@0.5, R1@0.7, and mAP for moment retrieval as in Lei et al. (2021); Moon et al. (2023); Lin et al. (2023). Highlight Detection is evaluated in terms of “Very Good” mAP and HiT@1. In Table 11(b), InternVideo2-1B and InternVideo2-6B bring gradual performance improvements compared to CLIP Radford et al. (2021) and CLIP Radford et al. (2021)+Slowfast Feichtenhofer et al. (2019). This suggests that a larger spatiotemporal model is more beneficial to short-term video semantic alignment capabilities.
@@ 5.3 Video-centric Dialogue and its Applications
Tab. 14 shows results on MVBench Li et al. (2023d), Egoschema Mangalam et al. (2023), and Perception Test Patraucean et al. (2024), by equipping VideoChat2 Li et al. (2023d) with our InternVideo2 encoder, compared with other MLLMs. Note VideoChat2-HD-F16 with InternVideo2 outperforms other systems with a clear margin in Perception Test and MVBench (F16 denotes the model is trained and tested with 16-frame inputs), except in Egoschema compared with proprietary commercial models (GPT4 & Gemini). Our method demonstrates superior short-term fine-grained video understanding compared to both open-source and commercial methods. Egoschema requires longer context harnessing, and we plan to further explore leveraging LLM for long multimodal sequence modeling. Considering these benchmarks not only involve perception but also reasoning, these results suggest InternVideo2 does embed knowledge to partially model the world. It also validates the importance of learning a transferrable video representation for current video-related MLLM. We also give several qualitative evaluations with popular GPT-4V and GeminiPro, including dialogues on action sequence (Fig. 4), confused action (Fig. 5), temporal order understanding (Fig. 6), temporal event counting (Fig. 7, unexpected action reasoning (Fig. 8), and vision-language navigation (Fig. 9).
@@ 5.4.2 Training Data and used Teachers in Stage 1
In Tab. 24, we examine the impact of distillation teachers and used dataset sizes on model performance. (a) Data Scale: Note that pretraining data scale for MAE should grow with the increasing model scale, otherwise the downstream performance would be saturated, such as K710 (0.66M videos) for ViT-L, K-Mash1.1M for ViT-1B, and K-Mash2M for ViT-6B. (b) Teacher: Tab. 24 reveals that the synergy between a multimodal teacher (e.g., CLIP Radford et al. (2021)) and a motion-aware teacher (e.g., MAE Tong et al. (2022)) markedly boosts performance, especially on SthSthV2. It highlights the importance of strategic teacher model selection in the distillation process.
@@ 5.4.4 Training and Evaluation in Stage 3
As shown in Tab. 19, incorporating questions (i.e., the ‘q’ in ‘QA’) into QFormer during Stage 3 training, which was found useful in Dai et al. (2023), actually harms the out-of-domain performance of the scaled-up VideoLLM. The NextQA training data is already included in the training corpus, and this is the only benchmark where the question-injected version performs better. Therefore, we believe that adding questions to QFormer during the instruction tuning stage of the scaled VideoChat model leads to some degree of overfitting.
@@ 6 Conclusion and Discussion
We have introduced a new family of video foundation models called InternVideo2, which achieves the state-of-the-art performance across various video and audio tasks. In InternVideo2, we combine masked video modeling, video-audio-text contrastive learning, and next token prediction into a unified framework. Additionally, we create a new video-text dataset that incorporates video-audio-speech fused captions as descriptions. The dataset contains temporally segmented clips with high semantic coherence. These designs in InternVideo2 contribute to enhancing video understanding in both perception and reasoning tasks. Notably, InternVideo2 excels in video-related dialogue and long video understanding, demonstrating its effectiveness in capturing high-level semantics.
@@
Despite its achievements, InternVideo2 does not introduce specific novel architectural design. Instead, it leverages the existing learning techniques for scaling video foundation models while focusing on improving data processing to enhance its spatiotemporal perception, semantic alignment, and basic knowledge embedding. Similarly to previous studies Li et al. (2023e); Ye et al. (2023), InternVideo2 still grapples with limitations stemming from fixed input resolutions, sampling rates, and highly compressed tokens, which restrict its ability to express rich video information and capture fine-grained details.
@@
Although InternVideo2 has demonstrated leading performance in some video understanding and reasoning benchmarks, it cannot guarantee an implicit world model that ensures consistency in visual reasoning. The inherent constraints imposed by fixed input representations, coupled with the complexity of visual reasoning tasks, present challenges in achieving a comprehensive and consistent understanding of the visual world.
@@
We investigate the potential biases here. We focus on age, gender, and race distributions, as these are commonly recognized areas where bias can occur. We count keywords related to these categories in the used captions. Note that these synthetic captions may not fully reflect the truth of the corresponding videos, thereby creating a gap between our analysis and the actual reality. Here are the results of our analysis: